# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 2.3 Suite of MAD simulations
We carry out simulations of MADs to compute the energy and angular momentum fluxes unto the BH. For this, we use H-AMR (Liska et al., 2022), a 3D massively parallel graphical processing unit (GPU) accelerated GRRMHD code, which is based on 3D HAMRPI (Tchekhovskoy, 2019) and 2D HARM2D (Gammie et al., 2003; Noble et al., 2006) codes. We initialize the simulations with a torus in hydrostatic equilibrium (Fishbone & Moncrief, 1976) about a BH, in a Kerr-Schild metric (see Sec. 2.3.12.3.3). We adopt spherical polar coordinates, $r$, $\theta$, $\varphi$, and use a uniform grid in $\log r$, $\theta$, and $\varphi$. The radial grid extends from just inside the event horizon, $R_{\rm in}=0.8r_{\rm h}$, to $R_{\rm out}=10^{5}r_{\rm g}$ (or $R_{\rm out}=10^{4}r_{\rm g}$). All our simulations have at least $5$ radial cells inside of the event horizon; this ensures that the exterior of the BH is causally disconnected from the inner radial boundary of the grid. We adopt outflow boundary conditions (BCs) at the inner and outer radial boundaries, transmissive BCs at the polar boundaries, and periodic BCs in the $\varphi$-direction (Liska et al., 2022).
@@
We insert a large poloidal magnetic flux loop into the torus, with the covariant vector potential described below. We wait for large-scale disk magnetic flux to flood the BH and lead to the MAD state (e.g., Tchekhovskoy et al., 2011).
@@
To compute the evolution of BH spin over cosmological timescales, which are far longer than the duration of any of our simulations, we perform a suite of GRMHD simulations spanning a wide range of BH spin values. We keep the BH spin constant throughout each of the simulations, a good approximation given the extremely short duration of our simulations in the cosmological context. However, in each of the simulations, we compute the torques acting on the BH. We then use these to model the BH spin evolution by solving a set of coupled ODEs. To understand how the disk thickness affects BH spin evolution, we have run the simulations with thermal disk scale heights ranging from thick, $h/r=0.3$, to thinner, $h/r=0.1$ and thinnest, $h/r=0.03$, disks. Figure 1 shows the vertical slices through density in simulations of MADs around rapidly spinning BH, $a=0.9375$, for these three disk thickness values. We see the that smaller $h/r$ disks are thinner and result in wider jets. Whereas most our simulations use a radiative cooling prescription (Noble et al., 2009), some use the full radiation transport scheme with a two-moment (M1) radiative closure (Liska et al., 2022). We summarize the main simulation parameters in Table 1 and discuss the simulations below in more detail.
@@
Table 1: Simulation details. Column 1: models. Column 2: BH spin. Column 3: thermal disk scale height. Column 4: polytropic index. Column 5: total spin-up parameter. Column 6: azimuthal wedge. Column 7: base grid resolution. Column 8: effective resolution within the disk. Column 9: simulation duration. Column 10: time-averaging window. Note ”*” indicates measurements on unresolved simulations performed for the resolution study described in Appendix A.
$a$ $h/r$ $\Gamma$ $s$ $\Delta_{\varphi}$ ($N_{r}\times N_{\theta}\times N_{\varphi}$)base ($N_{r}\times N_{\theta}\times N_{\varphi}$)eff $t_{\rm run}\,[r_{g}/c]$ $t_{\rm avg}\,[r_{g}/c]$
$\mathit{h/r=0.3}$
H3a-0.9 $-0.9$ $0.3$ $4/3$ 5.98 $2\pi$ $288\times 128\times 64$ $288\times 128\times 64$ ($0;20100$) ($10000;20100$)
H3a-0.5 $-0.5$ $0.3$ $4/3$ 5.33 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;16350$) ($10000;16350$)
H3a-0.2 $-0.2$ $0.3$ $4/3$ 3.61 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;15200$) ($10000;15200$)
H3a0.0 $0.0$ $0.3$ $4/3$ 1.03 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;18550$) ($10000;18550$)
H3a0.1 $0.1$ $0.3$ $4/3$0.5 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;18725$) ($10000;18725$)
H3a0.2 $0.2$ $0.3$ $4/3$1.56 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;13400$) ($10000;13400$)
H3a0.5 $0.5$ $0.3$ $4/3$4.62 $\pi$ $288\times 128\times 32$ $288\times 128\times 32$ ($0;13050$) ($10000;13050$)
H3a0.9 $0.9$ $0.3$ $4/3$7.47 $2\pi$ $288\times 128\times 64$ $288\times 128\times 64$ ($0;19900$) ($10000;19900$)
H3a0.99 $0.99$ $0.3$ $4/3$7.39 $2\pi$ $288\times 128\times 64$ $288\times 128\times 64$ ($0;14650$) ($10000;14650$)
$\mathit{h/r=0.3}$ $HR$
@@
H3a0.2hr 0.2 $0.3$ $13/9$ $-2.40$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;14800$) ($10000;14750$)
H3a0.5hr 0.5 $0.3$ $13/9$ $-5.31$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;57825$) ($52800;57800$)
H3a0.94hr 0.9375 $0.3$ $13/9$ $-8.67$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;68105$) ($63000;68000$)
@@
H1a0.3 0.3 $0.1$ $13/9$ $-0.09$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;58800$) ($53800;58800$)
H1a0.5 0.5 $0.1$ $13/9$ $-1.36$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;65600$) ($60600;65600$)
H1a0.94 $0.9375$ $0.1$ $13/9$ $-3.02$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;56300$) ($51300;66300$)
$\mathit{h/r=0.1}$ $HR$
H1a0.3hr 0.3 $0.1$ $13/9$ $-0.17$ $2\pi$ $512\times 288\times 256$ $1024\times 576\times 512$ ($0;59400$) ($45000;59400$)
$\mathit{h/r=0.05}$ $LR$ (warning: unresolved)
@@
H05a0.4 0.4 $0.05$ $13/9$ $-0.02^{*}$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;59380$) ($54400;59380$)
$\mathit{h/r=0.05}$ $HR$
H05a0.4hr 0.4 $0.05$ $13/9$ $-0.636$ $2\pi$ $512\times 288\times 256$ $1024\times 576\times 512$ ($0;45000$) ($41000;45000$)
@@
Ra0.4 0.4 $0.03$ $5/3$0.39 $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;190869$) ($188889;190869$)
Ra0.94 $0.9375$ $0.03$ $5/3$ $-3.18$ $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;192879$) ($188889;192879$)
@@
H03a0.94 $0.9375$ $0.03$ $5/3$ $-2.854$ $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;203339$) ($188839;203339$)
@@
Figure 1: Smaller $h/r$ disks (shown in yellow) are thinner and result in wider, less collimated polar jets (dark purple), as seen in the vertical slices through density for simulations with different disk scale heights, $h/r=0.3$ (model H3a0.94hr), $0.1$ (model H1a0.94), and $0.03$ (model Ra0.94). Black lines, which show axisymmetric poloidal magnetic flux contours, demonstrate that both the BH and inner disk are flooded by the vertical magnetic flux, as expected in MADs.
2.3.1 Nonradiative thick MADs
@@
To study thick MADs we utilize the simulations of Tchekhovskoy et al. (2011, 2012); Tchekhovskoy (2015) with BH spins $-0.9$, $-0.5$, $-0.2$, $0.0$, $0.1$, $0.2$, $0.5$, $0.9$, and $0.99$ and a polytropic index of $\Gamma=4/3$. These simulations are denoted as $h/r=0.3$ in Table 1. Collectively, we refer to these simulations as $h/r=0.3$ simulations. Compared to other simulations in this paper, the Tchekhovskoy et al. (2011) simulations are lower-resolution (at most $288\times 128\times 128$) and shorter-duration ($t_{\rm final}\leq 30,500r_{\rm g}/c$). Here, we have run higher-resolution ($384\times 300\times 64$) and longer simulations for five spin values, $a=-0.9$, $-0.5$, $0.2$, $0.5$, and $0.9375$, with a polytropic index of $\Gamma=13/9$. Collectively, we refer to these simulations as $h/r=0.3\,HR$. We show these higher-resolution thick disk MAD simulations in Table 1. Because our thick nonradiative MADs reach a thermal scale height of $h/r\approx 0.3$, we use this fiducial value to designate these simulations.
2.3.2 Cooled MADs
@@
For two sets of models, with $h/r=0.1$ and $h/r=0.05$, we cool the disk to its target thickness on the Keplerian timescale using a cooling source term following Noble et al. (2009). For the initial conditions, we adopt the torus of inner radius, $r_{\rm in}=6.5r_{\rm g}$, and the pressure maximum radius, $r_{\rm max}\sim 80r_{\rm g}$, where we tweak the exact value of $r_{\rm max}$ such that our torus extends out to an extremely large (but finite) distance, $r_{\rm out}\sim\text{few}\times 10^{4}r_{\rm g}$. To quickly achieve the MAD state, we adopt the covariant vector potential, $A_{\varphi}=q^{2}$, where (EHT Code Comparison Paper, in preparation)
$q=\frac{\rho}{\max\rho}\left(\frac{r}{r_{\rm in}}\right)^{3}\sin^{3}{\theta}\exp\left(-\frac{r}{400r_{\rm g}}\right)-0.2.$ (11)
We then normalize the magnetic field strength such that the minimum ratio of gas to magnetic pressure is $\min\beta=100$.
@@
We have carried out both low- and high-resolution simulations for $h/r=0.1$ and $h/r=0.05$. The low-resolution runs use a base grid of $384\times 300\times 64$ with no grid refinement, while the high-resolution runs use a base grid of $512\times 288\times 256$ with one level of static mesh refinement (SMR) applied for $|\theta-\pi/2|\leq 0.315$ and $4\,r_{g}\leq r\leq 500\,r_{g}$, resulting in an effective resolution of $1024\times 576\times 512$ in the disk proper. A resolution convergence study (see Appendix A) confirms that both resolutions are adequate for $h/r=0.1$ simulations. However, thinner disks ($h/r=0.05$) are only resolved at the high resolution, so only the high-resolution $h/r=0.05$ simulations are used in our analysis.
@@
Additionally, we include a very high-resolution $h/r=0.03$ simulation with Noble et al. (2009) cooling, previously presented in Liska et al. (2022). We refer to this simulation as model H03a0.94 and also label it as $h/r=0.03$ $1T$. Its resolution matches that of the radiative (GRRMHD) simulations (models RaX) described in Section 2.3.3. Its initial conditions are described in Liska et al. (2022).
2.3.3 Thin radiative MADs

### v2

@@ II.3 Suite of MAD simulations
We carry out simulations of MADs to compute the energy and angular momentum fluxes into the BH. For this, we use H-AMR [50], a 3D massively parallel graphical processing unit (GPU) accelerated GRRMHD code, which is based on 3D HAMRPI [78, 79, 97] and 2D HARM2D [24, 71] codes. We initialize the simulations with a torus in hydrostatic equilibrium [21] about a BH, in a Kerr-Schild metric (see Sec. II.3.1II.3.3). We adopt spherical polar coordinates, $r$, $\theta$, $\varphi$, and use a uniform grid in $\ln r$, $\theta$, and $\varphi$. The radial grid extends from just inside the event horizon, $R_{\mathrm{in}}=0.8r_{\mathrm{H}}$, to $R_{\mathrm{out}}=10^{5}r_{\mathrm{g}}$ (or $R_{\mathrm{out}}=10^{4}r_{\mathrm{g}}$). All our simulations have at least five radial cells inside of the event horizon; this ensures that the exterior of the BH is causally disconnected from the inner radial boundary of the grid. We adopt outflow boundary conditions (BCs) at the inner and outer radial boundaries, transmissive BCs at the polar boundaries, and periodic BCs in the $\varphi$-direction [50].
@@
We insert a large poloidal (i.e., with magnetic field pointing in the $R$- and $z$-directions) magnetic flux loop into the torus, with the covariant magnetic vector potential described below. We wait for the large-scale disk magnetic flux to flood the BH and lead to the MAD state [e.g., 95].
@@
To compute the evolution of the BH spin over cosmological timescales, which are far longer than the duration of any of our simulations, we perform a suite of GRMHD simulations spanning a wide range of BH spin values. We keep the BH spin constant throughout each of the simulations, a good approximation given the extremely short duration of our simulations in the cosmological context. However, in each of the simulations, we compute the torques acting on the BH. We then use these to model the BH spin evolution by solving a set of coupled ODEs. To understand how the disk thickness affects the BH spin evolution, we have run the simulations with dimensionless thermal disk scale heights ranging from thick, $h/r=0.3$, to thinner, $h/r=0.1$ and thinnest, $h/r=0.03$, disk.
Figure 1 shows the vertical slices through density in our simulations of MADs around a rapidly spinning BH, $a=0.9375$, for these three disk thickness values. We see that the smaller $h/r$ disks are thinner and result in wider jets. Whereas most our simulations use a radiative cooling prescription, for which $h=\sqrt{2/\pi}c_{\mathrm{i}}/\Omega$, where $c_{\mathrm{i}}$ is the isothermal sound speed and $\Omega$ is the angular frequency of the gas [72], some use the full radiation transport scheme with a two-moment (M1) radiative closure [51]. We summarize the main simulation parameters in Table 1 and discuss the simulations below in more detail.
@@
Table 1: Simulation details. Column 1: model name. Column 2: BH spin. Column 3: thermal disk scale height. Column 4: polytropic index. Column 5: spin-up parameter. Column 6: azimuthal wedge. Column 7: base grid resolution. Column 8: effective resolution within the disk. Column 9: simulation duration. Column 10: time-averaging window.
$a$ $h/r$ $\Gamma$ $s$ $\Delta_{\varphi}$ ($N_{r}\times N_{\theta}\times N_{\varphi}$)base ($N_{r}\times N_{\theta}\times N_{\varphi}$)eff $t_{\mathrm{run}}\,[r_{\mathrm{g}}/c]$ $t_{\mathrm{avg}}\,[r_{\mathrm{g}}/c]$
$\mathit{h/r=0.3}$, $\mathit{\Gamma=4/3}$
H3a-0.9 0.9-0.9 $0.3$ $4/3$ 5.98 $2\pi$ $8\times 128\times 64 288×2$$8\times 224\times 64$ ($0;20100$) ($10000;20100$)
H3a-0.5 0.5-0.5 $0.3$ $4/3$ 5.33 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;16350$) ($10000;16350$)
H3a-0.2 0.2-0.2 $0.3$ $4/3$ 3.61 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;15200$) ($10000;15200$)
H3a0.0 $0.0$ $0.3$ $4/3$ 1.03 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;18550$) ($10000;18550$)
H3a0.1 $0.1$ $0.3$ $4/3$ -0.5 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;18725$) ($10000;18725$)
H3a0.2 $0.2$ $0.3$ $4/3$ -1.56 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;13400$) ($10000;13400$)
H3a0.5 $0.5$ $0.3$ $4/3$ -4.62 $\pi$ $8\times 128\times 32 288×2$$8\times 224\times 32$ ($0;13050$) ($10000;13050$)
H3a0.9 $0.9$ $0.3$ $4/3$ -7.47 $2\pi$ $8\times 128\times 64 288×2$$8\times 224\times 64$ ($0;19900$) ($10000;19900$)
H3a0.99 $0.99$ $0.3$ $4/3$ -7.39 $2\pi$ $8\times 128\times 64 288×2$$8\times 224\times 64$ ($0;14650$) ($10000;14650$)
$\mathit{h/r=0.3}$ HRHR, $\mathit{\Gamma=13/9}$
@@
H3a0.2hr 0.2 $0.3$ $13/9$ 2.40-2.40 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;14800$) ($10000;14750$)
H3a0.5hr 0.5 $0.3$ $13/9$ 5.31-5.31 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;57825$) ($52800;57800$)
H3a0.94hr 0.9375 $0.3$ $13/9$ 8.67-8.67 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;68105$) ($63000;68000$)
@@
H1a0.3 0.3 $0.1$ $13/9$ 0.09-0.09 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;58800$) ($53800;58800$)
H1a0.5 0.5 $0.1$ $13/9$ 1.36-1.36 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;65600$) ($60600;65600$)
H1a0.94 $0.9375$ $0.1$ $13/9$ 3.02-3.02 $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;56300$) ($51300;66300$)
$\mathit{h/r=0.1}$ HRHR
H1a0.3hr 0.3 $0.1$ $13/9$ 0.17-0.17 $2\pi$ $512\times 288\times 256$ $1024\times 576\times 512$ ($0;59400$) ($45000;59400$)
$\mathit{h/r=0.05}$ LRLR (warning: unresolved)
@@
H05a0.4 0.4 $0.05$ $13/9$ $0.02^{*}$ $2\pi$ $384\times 300\times 64$ $384\times 300\times 64$ ($0;59380$) ($54400;59380$)
$\mathit{h/r=0.05}$ HRHR
H05a0.4hr 0.4 $0.05$ $13/9$ 0.636-0.636 $2\pi$ $512\times 288\times 256$ $1024\times 576\times 512$ ($0;45000$) ($41000;45000$)
@@
Ra0.4 0.4 $0.03$ $5/3$ -0.39 $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;190869$) ($188889;190869$)
Ra0.94 $0.9375$ $0.03$ $5/3$ 3.18-3.18 $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;192879$) ($188889;192879$)
@@
H03a0.94 $0.9375$ $0.03$ $5/3$ 2.854-2.854 $2\pi$ $1020\times 432\times 288$ $4080\times 1728\times 1152$ ($188839;203339$) ($188839;203339$)
[95] continued this run to later time ($t\sim 30,500r_{\mathrm{g}}/c$) at a higher resolution ($288\times 128\times 128$).
* Measurements on unresolved simulations performed for the resolution study described in Appendix A.
@@
Figure 1: Smaller $h/r$ MADs (yellow) are thinner and result in wider, less collimated polar jets (dark purple), as seen in the vertical slices through the instantaneous fluid-frame density for our simulations with different disk scale heights, $h/r=0.3$ (model H3a0.94hr), $0.1$ (model H1a0.94), and $0.03$ (model Ra0.94), as labeled on the panels. Black lines show the axisymmetric poloidal magnetic flux contours and demonstrate that the vertical magnetic flux floods both the BH and inner disk, as expected in MADs.
II.3.1 Nonradiative thick MADs
@@
To study thick MADs we utilize the simulations of [95, 92, 96] with BH spins a=0.9a=-0.9, 0.5-0.5, 0.2-0.2, $0.0$, $0.1$, $0.2$, $0.5$, $0.9$, and $0.99$ and a polytropic index of $\Gamma=4/3$. We label these simulations as $h/r=0.3$ in Table 1. Collectively, we refer to these simulations as the $h/r=0.3$ simulations. Compared to other simulations in this paper, these simulations are lower resolution (at most $288\times 128\times 128$) and shorter duration ($t_{\mathrm{final}}\leq 30,500r_{\mathrm{g}}/c$). Here, we have run higher poloidal resolution ($384\times 300\times 64$) and longer duration simulations for a range of spin values, a=0.9a=-0.9, 0.5-0.5, $0.2$, $0.5$, and $0.9375$, with a polytropic index of $\Gamma=13/9$. Collectively, we refer to these simulations as h/r=0.3HRh/r=0.3\,HR. We show these higher-resolution thick disk MAD simulations in Table 1. Because our thick nonradiative MADs reach a thermal scale height of $h/r\approx 0.3$, we use this fiducial value when referring to these simulations.
II.3.2 Thin MADs
@@
For two sets of our models, with $h/r=0.1$ and $h/r=0.05$, we cool the disk to its target thickness on the Keplerian timescale using a cooling source term following [72]. For the initial conditions, we adopt an equilibrium hydrodynamic torus [21] of inner radius, $r_{\mathrm{in}}=20r_{\mathrm{g}}$, and the pressure maximum radius, $r_{\mathrm{max}}\sim 41r_{\mathrm{g}}$, where we tweak the exact value of $r_{\mathrm{max}}$ such that our torus extends out to an extremely large (but finite) distance, $r_{\mathrm{out}}\sim\text{few}\times 10^{4}r_{\mathrm{g}}$. To quickly achieve the MAD state, we adopt the covariant vector potential, $A_{\varphi}=q^{2}$, where (EHT Code Comparison Paper, in preparation),
$q=\frac{\rho}{\max\rho}\left(\frac{r}{r_{\mathrm{in}}}\right)^{3}{\sin^{3}}\theta\,\exp\left(-\frac{r}{400r_{\mathrm{g}}}\right)-0.2,$ (11)
which results in a radially extended poloidal magnetic field loop of a sufficiently large poloidal magnetic flux to flood the BH. [^1] In Eq. (11), the $r^{3}$ term increases the size and amount of the magnetic flux in the loop by pushing the center of the loop out to larger radii. However, this results in a magnetically dominated outer disk. The exponential term reduces the outer disk magnetization by redistributing some of the magnetic flux back toward smaller radii. The last term in Eq. (11) ensures that the magnetic loop does not extend too close to the torus boundaries. We then normalize the magnetic field strength such that the minimum ratio of gas to magnetic pressure is $\min\beta=100$.
@@
We have carried out both low- and high-resolution simulations for $h/r=0.1$ and $h/r=0.05$. The low-resolution runs use a base grid of $384\times 300\times 64$ with no grid refinement, while the high-resolution runs use a base grid of $512\times 288\times 256$ with one level of static mesh refinement (SMR) applied for $|\theta-\pi/2|\leq 0.315$ and $4\,r_{\mathrm{g}}\leq r\leq 500\,r_{\mathrm{g}}$, resulting in an effective resolution of $1024\times 576\times 512$ in the disk proper. A resolution convergence study (see Appendix A) confirms that both resolutions are adequate for $h/r=0.1$ simulations. However, thinner disks ($h/r=0.05$) are only resolved at the high-resolution, so we use only the high-resolution $h/r=0.05$ simulations in our analysis.
@@
Additionally, we include a very high-resolution $h/r=0.03$ simulation using a predefined cooling function [72], previously presented in [51]. We refer to this simulation as model H03a0.94 and also label it as $h/r=0.03$ $1T$. Its resolution matches that of the radiative (GRRMHD) simulations (models RaX) described in Section II.3.3. Its initial conditions are described in [51].
II.3.3 Thin radiative MADs
Although cooled GRMHD models with fixed $h/r$ offer computational control, they prescribe thermodynamics rather than predict it. GRRMHD models, in contrast, include radiation transport and allow the disk structure to emerge self-consistently, providing the most physically realistic test of whether predefined cooling functions adequately capture the structure of radiatively cooled accretion disks.